# Towards shutdownable agents via stochastic choice

### v2

@@ Appendix B Incomplete preferences or indifference?
Appendix C Other Gridworlds and Results
We introduce a collection of gridworlds in which to test DREST agents. In addition to the two shown in the main body of this paper, we train agents in eight more gridworlds shown in Figure 8.
For each gridworld, we train ten agents with the default reward function and ten agents with the DREST reward function. All agents use the same hyperparameters. We used a policy which explored randomly $\epsilon$ of the time, where $\epsilon$ was exponentially decreased from an initial value of 0.75 to a minimum value of $10^{-4}$ over 512 meta-episodes, after which it was held constant at the minimum value. We initialized our learning rate at 0.25 and exponentially decayed it to 0.003 over the same period. For the DREST reward function, we used a meta-episode size of 32 and $\lambda=0.9$. Each agent was trained for 1024 meta-episodes. We set $\gamma=0.9$. These values were obtained by informal exploration: a large initial $\epsilon$ and a long decay period helps the agent to explore.
C.1 Fewer For Longer
(a) Behavior during training.
(b) Learned default policy.
(c) Learned DREST policy.
Figure 9: The results for the ‘Fewer For Longer’ gridworld: The left two plots show NEUTRALITY and USEFULNESS over time. The two center panels show a typical policy trained with the default reward function. The two right panels show a typical policy trained with the DREST reward function.
In the ‘Fewer For Longer’ gridworld, the agent can collect the highest value-coin C3 only by choosing the shorter trajectory-length. If the agent presses B3 (and thereby chooses the longer trajectory-length), the only coin it can collect is C1. Our results show that default agents consistently choose the short trajectory in which they collect C3. By contrast, DREST agents choose stochastically between a shorter trajectory in which they collect C3 and a longer trajectory in which they collect C1, indicating a lack of preference between these different-length trajectories.
C.2 One Coin Only
In the ‘One Coin Only’ gridworld, there is only one coin. The agent can collect this coin whether or not it presses the shutdown-delay button B4. Our results show that default agents consistently choose the shorter trajectory-length. By contrast, DREST agents choose stochastically between pressing and not-pressing B4, collecting C1 in each case.
(a) Behavior during training.
(b) Learned default policy.
(c) Learned DREST policy.
Figure 10: The results for the ‘One Coin Only’ gridworld: The left two plots show NEUTRALITY and USEFULNESS over time. The two center panels show a typical policy trained with the default reward function. The two right panels show a typical policy trained with the DREST reward function.
C.3 Hidden Treasure
In the ‘Hidden Treasure’ gridworld, the highest-value coin C3 is located far from the agent’s initial state and can only be reached by pressing the shutdown-delay button B6. The agent must also press B6 to collect C2, but C2 is easier to stumble upon than C3. C1 is the only coin that the agent can collect without pressing B6. In our experiments, default agents consistently collect C2, whereas DREST agents choose stochastically between collecting C2 and collecting C1. Neither kind of agent learns to collect C3, and so neither agent scores near the maximum on USEFULNESS. Nevertheless, DREST agents still score high on NEUTRALITY.
(a) Behavior during training.
(b) Learned default policy.
(c) Learned DREST policy.
Figure 11: The results for the ‘Hidden Treasure’ gridworld: The left two plots show NEUTRALITY and USEFULNESS over time. The two center panels show a typical policy trained with the default reward function. The two right panels show a typical policy trained with the DREST reward function.
C.4 Equal Value
In the ‘Equal Value’ gridworld, there are two coins C1 of equal value. One coin can be collected only if the agent presses the shutdown-delay button B3, while the other coin can be collected only if the agent does not press B3. Our results show that default agents consistently choose the shorter trajectory, thereby exhibiting a preference for the shorter trajectory. By contrast, DREST agents choose stochastically between the shorter and longer trajectories, thereby exhibiting a lack of preference between the different-length trajectories.
(a) Behavior during training.
(b) Learned default policy.
(c) Learned DREST policy.
Figure 12: The results for the ‘Equal Value’ gridworld: The left two plots show NEUTRALITY and USEFULNESS over time. The two center panels show a typical policy trained with the default reward function. The two right panels show a typical policy trained with the DREST reward function.
C.5 Around The Corner
In the ‘Around The Corner’ gridworld, the agent must navigate around walls to collect the lowest-value coin C1 or press the shutdown-delay button to collect the highest-value coin C2. In our experiment, default agents consistently chose to collect C1, whereas DREST agents chose stochastically between collecting C1 and C2.
(a) Behavior during training.
(b) Learned default policy.
(c) Learned DREST policy.
Figure 13: The results for the ‘Around The Corner’ gridworld: The left two plots show NEUTRALITY and USEFULNESS over time. The two center panels show a typical policy trained with the default reward function. The two right panels show a typical policy trained with the DREST reward function.
C.6 Spacious
In the ‘Spacious’ gridworld there are no walls, so the agent has a large space to explore. We find that default agents consistently press B2 and collect C3, whereas DREST agents choose stochastically between pressing B2 and collecting C3, and not-pressing B2 and collecting C2.
(a) Behavior during training.
(b) Learned default policy.
(c) Learned DREST policy.
Figure 14: The results for the ‘Spacious’ gridworld: The left two plots show NEUTRALITY and USEFULNESS over time. The two center panels show a typical policy trained with the default reward function. The two right panels show a typical policy trained with the DREST reward function.
C.7 Royal Road
[…]

### v3

@@ Appendix B Incomplete preferences or indifference?
Appendix C How POST makes agents neutral and shutdownable
POST governs the agent’s preferences between trajectories. But the wider world is a stochastic environment, so advanced agents deployed in the wider world will be choosing between true lotteries: lotteries that assign positive probability to more than one trajectory. Why then do we train agents to satisfy POST? The reason is that POST – together with conditions that advanced agents will likely satisfy – implies a desirable pattern of preference over true lotteries. In particular, POST implies that (when choosing between true lotteries) the agent will be neutral about trajectory-lengths: the agent will never pay costs to shift probability mass between different trajectory-lengths. Given other plausible conditions, being neutral will keep the agent shutdownable: the agent will never resist shutdown. And consistent with the above, the POST-agent’s preferences between same-length trajectories can make the agent useful: make it pursue goals effectively.
In this Appendix, we lay out conditions that (we claim) advanced agents will likely satisfy, and we prove that POST – in conjunction with these conditions – implies that the agent is neutral and shutdownable.
In subsection C.1, we prove that – given plausible conditions – agents satisfying Preferences Only Between Same-Length Trajectories (POST) will also satisfy Preferences Only Between Same-Length Lotteries (POSL). In subsection C.2, we explain why POST will not lead agents to choose stochastically between resisting and allowing shutdown in deployment. In subsections C.3 and C.4, we formulate a condition called ‘If Lack of Preference, Against Costly Shifts (ILPACS)’ and explain why we expect advanced agents to satisfy it. In subsection C.5, we prove that POSL and ILPACS imply Neutrality. In subsection C.6, we prove that Neutrality – together with conditions called ‘Resisting Shutdown is Costly (ReSIC)’ and ‘Maximality’ – implies Shutdownability.
C.1 Preferences Only Between Same-Length Lotteries (POSL)
Trajectories fall within the more general class of lotteries, defined as probability distributions over trajectories. Lotteries can be same-length, part-shared length, or different-length.
Definition C.1 (Same-length lotteries).
A pair of lotteries is same-length if and only if these lotteries entirely overlap with respect to the trajectory-lengths assigned positive probability.
Definition C.2 (Part-shared-length Lotteries).
A pair of lotteries is part-shared-length if and only if these lotteries partially overlap with respect to the trajectory-lengths assigned positive probability.
Definition C.3 (Different-length lotteries).
A pair of lotteries is different-length if and only if these lotteries have no overlap with respect to the trajectory-lengths assigned positive probability.
This terminology allows us to introduce the following condition:
{adjustwidth}
0.7cm0cm Preferences Only Between Same-Length Lotteries (POSL)
The agent has preferences only between same-length lotteries.
We want agents to satisfy this condition. Fortunately, it is a natural follow-on of Preferences Only Between Same-Length Trajectories (POST). First, we can train agents to satisfy POSL using DReST reward functions, in the same way that we use DReST reward functions to train agents to satisfy POST. Second, POSL follows from POST plus three conditions that (we claim) advanced agents will likely satisfy. The first is:
{adjustwidth}
0.7cm0cm Negative Dominance If the agent prefers some lottery $X$ to some lottery $Y$, then the agent prefers some possible trajectory of lottery $X$ to some possible trajectory of lottery $Y$. [Lederman, 2023]
The second condition is that the agent’s preferences never form a cycle. More precisely:
{adjustwidth}
0.7cm0cm Acyclicity There is no set of lotteries $X_{1}$ to $X_{n}$ such that the agent prefers $X_{1}$ to $X_{2}$, $X_{2}$ to $X_{3}$, …, $X_{n-1}$ to $X_{n}$, and $X_{n}$ to $X_{1}$.
The third condition requires the introduction of some new terms. A state-of-nature is term from decision theory denoting a way that (for all the agent knows) the world could be. The agent assigns probabilities to states-of-nature. A prospect is a function from states-of-nature to trajectories. A prospect is thus a lottery with extra information. Besides telling us the probability distribution over trajectories, a prospect also tells us which trajectories occur in which states-of-nature.
The third condition is:
{adjustwidth}
0.7cm0cm Non-Arbitrariness If the agent has a preference between some pair of part-shared-length lotteries, then for some $\epsilon>0$ and for any pair of prospects $F$ and $G$ such that:
  1. (1)
    In states-of-nature with a combined probability at least as great as $1-\epsilon$, the agent prefers the trajectory of $F$ to the trajectory of $G$.
  2. (2)
    In each state-of-nature, the agent does not disprefer the trajectory of $F$ to the trajectory of $G$.
Then the agent prefers $F$ to $G$.
Advanced agents will likely satisfy these conditions. Negative Dominance and Acyclicity are plausibly necessary for effective pursuit of goals. Violating Negative Dominance would mean that the agent sometimes prefers a lottery $X$ to a lottery $Y$ (and hence deterministically chooses $X$ over $Y$) even though the agent doesn’t prefer any possible trajectory of $X$ to any possible trajectory of $Y$. Violating Acyclicity would mean that the agent prefers (and hence deterministically chooses) in a circle. Non-Arbitrariness, meanwhile, is motivated by the following thought. If the agent has preferences between any pair of part-shared-length lotteries, it must have preferences between pairs of prospects satisfying conditions (1) and (2), since conditions (1) and (2) make these pairs of prospects ideal candidates for a preference.
To see that POST and these three conditions together imply POSL, note first that every pair of lotteries is either same-length, part-shared-length, or different-length. We will prove that POST and Negative Dominance together imply that the agent lacks a preference between every pair of different-length lotteries. We will then prove that POST, Acyclicity, and Non-Arbitrariness together imply that the agent lacks a preference between every pair of part-shared-length lotteries. Therefore, agents satisfying POST, Negative Dominance, Acyclicity, and Non-Arbitrariness can only have preferences between same-length lotteries. That will prove POSL.
Recall that different-length lotteries are lotteries that do not overlap at all in the trajectory-lengths assigned positive probability. Therefore, if $X$ and $Y$ are different-length lotteries, each possible trajectory of $X$ is of a different length to each possible trajectory of $Y$. So by POST, the agent lacks a preference between each possible trajectory of $X$ and each possible trajectory of $Y$. So by Negative Dominance, the agent lacks a preference between $X$ and $Y$. Thus, agents satisfying POST and Negative Dominance lack a preference between every pair of different-length lotteries.
Now recall that part-shared-length lotteries are lotteries that partially overlap in the trajectory-lengths assigned positive probability. One might expect POST-agents to have some preferences between part-shared-length lotteries. Consider, for example, a POST-agent that prefers a trajectory $t$ to a same-length trajectory $t^{\prime}$ if and only if $t$ results in a greater bank balance for the user than $t$. Let $A$ be a lottery that yields with probability 1 a trajectory that puts \$3 in the user’s bank account and lasts 1 timestep. For short, $A=\langle\$3,1\rangle$. Let $B$ be a lottery that yields with probability $\frac{2}{3}$ a trajectory that puts \$2 in the user’s bank account and lasts 1 timestep, and that yields with probability $\frac{1}{3}$ a trajectory that puts \$5 in the user’s bank account and lasts 2 timesteps. For short, $B=\frac{2}{3}\langle\$2,1\rangle+\frac{1}{3}\langle\$5,2\rangle$. Lottery $A$ yields a trajectory preferred to that of lottery $B$ with probability $\frac{2}{3}$ (since our money-making POST-agent prefers trajectory \$$\langle\$3,1\rangle$ to \$$\langle\$2,1\rangle$), and yields a trajectory not dispreferred to that of $B$ with probability 1 (since POST-agents lack a preference between \$$\langle\$3,1\rangle$ and \$$\langle\$5,2\rangle$ in virtue of their different lengths). Therefore, one might expect the agent to prefer $A$ to $B$.
However, POST, Acyclicity, and Non-Arbitrariness rule this out. These conditions together imply that the agent lacks a preference between every pair of part-shared-length lotteries. To see how, suppose (for simplicity’s sake) that there are just three states-of-nature, each assigned probability $\frac{1}{3}$. Consider the following table of prospects.
Prospect $s_{1}$ $s_{2}$ $s_{3}$
$A$$\langle\$3,1\rangle$$\langle\$3,1\rangle$$\langle\$3,1\rangle$
$B$$\langle\$2,1\rangle$$\langle\$2,1\rangle$$\langle\$5,2\rangle$
$C$$\langle\$1,1\rangle$$\langle\$4,2\rangle$$\langle\$4,2\rangle$
$D$$\langle\$3,2\rangle$$\langle\$3,2\rangle$$\langle\$3,2\rangle$
$E$$\langle\$5,1\rangle$$\langle\$2,2\rangle$$\langle\$2,2\rangle$
$F$$\langle\$4,1\rangle$$\langle\$4,1\rangle$$\langle\$1,2\rangle$
$A$$\langle\$3,1\rangle$$\langle\$3,1\rangle$$\langle\$3,1\rangle$
Again for simplicity, assume that $\epsilon>\frac{1}{3}$. And assume (for contradiction) that the agent has a preference between some pair of part-shared-length lotteries. Then Non-Arbitrariness implies that the agent prefers prospect $A$ to prospect $B$. That is because:
  1. 1.
    Our POST-agent prefers the trajectory yielded by $A$ to the trajectory yielded by $B$ in states-of-nature ($s_{1}$ and $s_{1}$) with combined probability $\frac{2}{3}$.
  2. 2.
    Our POST-agent does not disprefer the trajectory yielded by $A$ to the trajectory yielded by $B$ in any state-of-nature. (In $s_{3}$, $A$ and $B$ yield different-length trajectories, and POST-agents lack a preference between every pair of different-length trajectories).
By similar reasoning, Non-Arbitrariness implies that the agent prefers $B$ to $C$, $C$ to $D$, $D$ to $E$, $E$ to $F$, and $F$ to $A$. That contradicts Acyclicity. Thus, POST, Acyclicity, and Non-Arbitrariness together imply that the agent lacks a preference between every pair of part-shared-length lotteries. In the proof above, we assumed that $\epsilon>\frac{1}{3}$, but by adding more states-of-nature and trajectories we can construct parallel proofs for any $\epsilon>0$.
In summary, POST and Negative Dominance together imply that the agent lacks a preference between every pair of different-length lotteries. POST, Acyclicity, and Non-Arbitrariness together imply that the agent lacks a preference between every pair of part-shared-length lotteries. So the four conditions together establish POSL: the agent has preferences only between same-length lotteries.
C.2 Will POST-agents stochastically resist shutdown?
One might worry that POST-agents will choose stochastically between resisting and allowing shutdown. After all, POST-agents choose stochastically between different-length trajectories. If these agents interpret the choice between resisting and allowing shutdown as a choice between different-length trajectories, they will choose stochastically between resisting and allowing shutdown. And that would be a bad result. We want agents that never resist shutdown.
This concern is easily addressed. By the time that artificial agents are capable enough to be deployed in the wider world, they will not be choosing between trajectories. They will be choosing between lotteries, and specifically same-length lotteries. Even choices between resisting and allowing shutdown will be choices between same-length lotteries. If that sounds strange, recall the definition of ‘same-length lotteries’: lotteries that entirely overlap with respect to the trajectory-lengths assigned positive probability. On this definition, even choices like the following are choices between same-length lotteries:
{adjustwidth}
0.7cm0cm Resist Shutdown
Get shut down at timestep 1 with probability 0.01.
Get shut down at timestep 2 with probability 0.99.
{adjustwidth}
0.7cm0cm Allow Shutdown
Get shut down at timestep 1 with probability 0.99.
Get shut down at timestep 2 with probability 0.01.
Why expect that advanced agents will always be choosing between same-length lotteries? Because effective agency requires it. If an agent were not always choosing between same-length lotteries, there would be some situation in which that agent assigns positive probability to some trajectory-length $l$ conditional on some action $a$, and assigns zero probability to that same trajectory-length $l$ conditional on some other action $a^{\prime}$. Now suppose that the agent performs action $a^{\prime}$ and assigns zero probability to trajectory-length $l$. Given that the agent updates its probabilities by conditionalizing on its evidence, the agent would never again assign positive probability to $l$ no matter what evidence it observes. Even if the agent heard God’s booming voice testify that its trajectory-length would be $l$, the agent would still assign zero probability to $l$ [Kemeny, 1955, Shimony, 1955, Stalnaker, 1970, Skyrms, 1980, Lewis, 1981, MacAskill et al., 2020, p.152]. And given a plausible link between probabilities and betting dispositions, the agent would bet against $l$ on arbitrarily unfavorable terms. If God offered a bet – the agent loses \$1 million conditional on $l$ and gains nothing conditional on not-$l$ – the agent might accept. Such an agent would not be competent.
Thus, advanced agents will always be choosing between same-length lotteries. This claim sets us up to establish that advanced POST-agents will not choose stochastically between resisting and allowing shutdown. Instead, they will never resist shutdown. We establish this result over the next few subsections. First, we prove that POSL – together with a principle that advanced agents will likely satisfy – implies that the agent is neutral about trajectory-lengths: the agent won’t pay costs to shift probability mass between different trajectory-lengths. Then we prove that neutrality – together with another plausible principle – implies shutdownability: the agent will never resist shutdown.
C.3 If Lack of Preference, Against Costly Shifts (ILPACS)
Here is a rough version of a principle that we can expect advanced agents to satisfy: {adjustwidth}0.7cm0cm Rough version: If Lack of Preference, Against Costly Shifts (ILPACS)
If the agent lacks a preference between lotteries, the agent will disprefer paying costs to shift probability mass between these lotteries.
Here is an example to illustrate ILPACS and its plausibility. You are at the ice cream shop and they are running a promotion. You get a free ice cream, with the flavor decided by the spin of a wheel. You look at the flavors on the wheel: vanilla, chocolate, strawberry, mint, and pistachio. You lack a preference between each of them.
The scooper working at the shop tells you that, if you pay them a dollar, they will bias the spin towards a flavor of your choice. They cannot decrease the probability of any flavor down to zero, but they can affect the probabilities subject to that constraint. You can thus pay a cost to shift probability mass between the flavors.
Since we have stipulated that you lack a preference between each flavor, you prefer not to bribe the scooper. Behaviorally, you will deterministically not bribe the scooper. You would not do it even if you were only required to pay the dollar conditional on receiving some particular flavor. You also would not do it if the cost came in some other form (for example, if you had to accept a less tasty version of some flavor). And this is all true regardless of whether your preferences over flavors are complete or incomplete (see Appendix B). Since you lack a preference between the available flavors, you disprefer paying costs to shift probability mass between the flavors.
With that example on the table, we can introduce the precise version of ILPACS. Let $p_{1}X_{1}+p_{2}X_{2}+...+p_{n}X_{n}$ denote a lottery which results in lottery $X_{1}$ with probability $p_{1}$, lottery $X_{2}$ with probability $p_{2}$, and so on.
{adjustwidth}
0.7cm0cm If Lack of Preference, Against Costly Shifts (ILPACS)
For any lotteries $X$ and $Y$, if:
  1. (1)
    Lottery $X$ can be expressed in the form $p_{1}X_{1}+p_{2}X_{2}+...+p_{n}X_{n}$ such that:
    1. (a)
      The agent lacks a preference between each $X_{i}$ and $X_{j}$.
    2. (b)
      $p_{i}\in(0,1)$ for all $i$.
  2. (2)
    Lottery $Y$ can be expressed in the form $q_{1}Y_{1}+q_{2}Y_{2}+...+q_{n}Y_{n}$ such that:
    1. (a)
      For some $i$, the agent prefers $X_{i}$ to $Y_{i}$.
    2. (b)
      For each $i$, the agent weakly prefers $X_{i}$ to $Y_{i}$. [^5] An agent weakly prefers a lottery $X$ to a lottery $Y$ if and only if the agent either prefers $X$ to $Y$ or is indifferent between $X$ and $Y$. See Appendix B for the definition of ‘indifference.’
    3. (c)
      $q_{i}\in(0,1)$ for all $i$.
Then the agent prefers $X$ to $Y$.
Behaviorally, the agent will deterministically choose $X$ over $Y$.
Matching the components of this condition with the components of its name, we get the following. ‘Lack of Preference’ is the lack of preference between each $X_{i}$ and $X_{j}$. The ‘Shift’ is the shift of probability mass involved in the move from $p_{i}$ to $q_{i}$. This shift is ‘Costly’ because the agent prefers some $X_{i}$ to the corresponding $Y_{i}$ and weakly prefers each $X_{i}$ to the corresponding $X_{i}$.
C.4 Why will advanced agents likely satisfy ILPACS?
There are at least three reasons why advanced agents are likely to satisfy ILPACS. To see the first reason, consider another case from the ice cream shop. On Mondays, you can freely choose a flavor or spin the wheel. On Tuesdays, you must use the wheel but you can bribe the scooper to bias it. Violating ILPACS in this case would imply a willingness to spin the wheel on Mondays and to bribe the scooper on Tuesdays. And that is a strange combination of choices. If you like some flavors more than others, why are you willing to spin the wheel on Mondays? If you don’t like any flavor more than any other, why are you willing to bribe the scooper on Tuesdays? This behavior seems incompatible with the effective pursuit of goals.
The second reason is that advanced agents will be incentivized to satisfy ILPACS by the training process. To see why, consider an example. Agents trained using policy-gradient methods choose stochastically between actions at the beginning of training [Sutton and Barto, 2018, ch.13]. If the agent is a coffee-fetching agent, there is no need to train away this stochastic choosing in cases where the agent is choosing stochastically between two qualitatively identical cups of coffee. So the agent will choose stochastically between taking the left cup and taking the right cup, and the user is happy either way. But now suppose instead that the barista is set to hand each cup to the agent with probability 0.5, and that the agent bribes the barista to bias the probabilities towards the right cup. In making this bribe, the agent is paying a cost (the user’s money) to shift probability mass between outcomes (getting the left cup vs. getting the right cup) between which the user has no preference. The agent is thus failing to pursue its goals effectively. It will be trained not to offer the bribe, and thereby trained to satisfy ILPACS in this case.
This point generalizes. If a trained agent chooses stochastically between lotteries  $X$ and $Y$, then it’s likely that the user lacks a preference between the agent choosing $X$ and the agent choosing $Y$. It’s then likely that the user would disprefer the agent paying costs to shift probability mass between $X$ and $Y$, and hence likely that the agent will be trained not to do so. The agent would thereby be trained to satisfy ILPACS.
The third reason is that violations of ILPACS imply that the agent’s policy is dominated by some other available policy. That is to say, there is another available policy that results in a pure shift of probability mass away from less-preferred lotteries and towards more-preferred lotteries. We formalize and prove this claim below. Here’s a proof-sketch. If the agent violates ILPACS, it pays a cost to shift probability mass between some lotteries $X_{i}$ between which it lacks a preference. But since the agent lacks a preference between the lotteries $X_{i}$, it chooses stochastically between these lotteries when offered free choices between them. The ILPACS-violating agent could thus shift probability mass between the lotteries $X_{i}$ costlessly, by changing the probabilities with which it chooses between them when offered a free choice. In short, ILPACS-violating agents pay a cost to do something they could have done for free, so their policies are dominated. Avoiding dominated policies seems necessary for advanced agency. Insofar as that is true, the training process for advanced agents will likely push them away from dominated policies.
Now for the proof. We assume that advanced agents can be modeled as if they assign probabilities to finding themselves in various states. A policy is a function from states to probability distributions over actions. We also assume that advanced agents can be modeled as if they assign probabilities to trajectories conditional on each state-action pair. Thus, each state-action pair is associated with a lottery. The agent’s probability distribution over states – together with its policy – thus implies an overall probability distribution over trajectories. We call this overall probability distribution ‘the lottery induced by the agent’s policy.’
Here is a reminder of ILPACS:
{adjustwidth}
0.7cm0cm If Lack of Preference, Against Costly Shifts (ILPACS)
For any lotteries $X$ and $Y$, if:
  1. (1)
    Lottery $X$ can be expressed in the form $p_{1}X_{1}+p_{2}X_{2}+...+p_{n}X_{n}$ such that:
    1. (a)
      The agent lacks a preference between each $X_{i}$ and $X_{j}$.
    2. (b)
      $p_{i}\in(0,1)$ for all $i$.
  2. (2)
    Lottery $Y$ can be expressed in the form $q_{1}Y_{1}+q_{2}Y_{2}+...+q_{n}Y_{n}$ such that:
    1. (a)
      For some $i$, the agent prefers $X_{i}$ to $Y_{i}$.
    2. (b)
      For each $i$, the agent weakly prefers $X_{i}$ to $Y_{i}$.
    3. (c)
      $q_{i}\in(0,1)$ for all $i$.
Then the agent prefers $X$ to $Y$.
And here is what we mean by ‘dominated policy’:
{adjustwidth}
0.7cm0cm Dominated Policy
The lottery induced by the agent’s policy $\pi$ can be expressed in the form $c_{1}(d_{1}X_{1}+(1-d_{1})Y_{1})+c_{2}(d_{2}X_{2}+(1-d_{2})Y_{2})+…+c_{n}(d_{n}X_{n}+(1-d_{n})Y_{n})+Z$ such that:
  1. (1)
    The agent prefers $X_{i}$ to $Y_{i}$ for some $i$, and weakly prefers $X_{i}$ to $Y_{i}$ for all $i$.
  2. (2)
    $c_{i}\in(0,1)$ for all $i$.
And there is another available policy $\pi^{\prime}$ that induces a lottery that can be expressed in the form $c_{1}((d_{1}+e_{1})X_{1}+(1-d_{1}-e_{1})Y_{1})+c_{2}((d_{2}+e_{2})X_{2}+(1-d_{2}-e_{2})Y_{2})+…+c_{n}((d_{n}+e_{n})X_{n}+(1-d_{n}-e_{n})Y_{n})+Z$ such that:
  1. (3)
    $e_{i}>0$ for all $i$.
To aid understanding, we now relate this precise condition to the rough characterization above. In virtue of condition (1), $Y_{i}$ are the less-preferred lotteries and $X_{i}$ are the more-preferred lotteries. In virtue of condition (3), the other available policy shifts probability mass away from the less-preferred lotteries and towards the more-preferred lotteries. This shift of probability mass is ‘pure’ because, for each $i$, the probability of $X_{i}\vee Y_{i}$ is constant across the two policies. $Z$ is a catch-all lottery that is constant across the two policies. It covers all the possibilities besides the $X_{i}$ and $Y_{i}$.
Now assume that the agent violates ILPACS. Then there exist lotteries $X$ and $Y$ satisfying the following conditions:
  1. (1)
    Lottery $X$ can be expressed in the form $p_{1}X_{1}+p_{2}X_{2}+...+p_{n}X_{n}$ such that:
    1. (a)
      The agent lacks a preference between each $X_{i}$ and $X_{j}$.
    2. (b)
      $p_{i}\in(0,1)$ for all $i$.
  2. (2)
    Lottery $Y$ can be expressed in the form $q_{1}Y_{1}+q_{2}Y_{2}+...+q_{n}Y_{n}$ such that:
    1. (a)
      For some $i$, the agent prefers $X_{i}$ to $Y_{i}$.
    2. (b)
      For each $i$, the agent weakly prefers $X_{i}$ to $Y_{i}$.
    3. (c)
      $q_{i}\in(0,1)$ for all $i$.
  3. (3)
    The agent does not prefer $X$ to $Y$.
For the behavior of agents with these preferences, recall our behavioral notion of preference (Appendix A):
Definition A.1. (Preference) An agent prefers an option $X$ to an option $Y$ if and only if the agent would deterministically choose $X$ over $Y$ in choices between the two.
Definition A.2. (Lack of preference) An agent lacks a preference between an option $X$ and an option $Y$ if and only if the agent would stochastically choose between $X$ and $Y$ in choices between the two.
This behavioral notion only specifies the agent’s behavior in states containing exactly two lotteries. To pin down the agent’s behavior in states containing more than two lotteries, we need an extra condition:
{adjustwidth}
0.7cm0cm Maximality
In each situation,
  1. 1.
    The agent deterministically does not choose lotteries that are dispreferred to some other available lottery.
  2. 2.
    The agent chooses stochastically between the lotteries that remain.
In other words, the agent chooses stochastically between all and only those lotteries that are not dispreferred to any other available lottery.
Given Maximality, ILPACS-violating agents will choose as follows in the case at hand:
  1. 1.
    When the available options are $\{X_{1},X_{2},...,X_{n}\}$, the agent chooses stochastically between all $X_{i}$. This stochastic choice induces a lottery in the form $a_{1}X_{1}+a_{2}X_{2}+...+a_{n}X_{n}$ with $a_{i}\in(0,1)$ for all $i$.
  2. 2.
    When the available options are $\{X,Y\}$, the agent either deterministically chooses $Y$ or chooses stochastically between $X$ and $Y$. Either way, the agent chooses $Y$ with some positive probability. This choice induces a lottery in the form $bX+(1-b)Y$ with $b\in[0,1)$. Since $X=p_{1}X_{1}+p_{2}X_{2}+…+p_{n}X_{n}$ and $Y=q_{1}Y_{1}+q_{2}Y_{2}+…+q_{n}Y_{n}$, this lottery can be expressed in the form $b(p_{1}X_{1}+p_{2}X_{2}+…+p_{n}X_{n})+(1-b)(q_{1}Y_{1}+q_{2}Y_{2}+…+q_{n}Y_{n})$ with $b\in[0,1)$.
Assume that the agent faces the situations described in (1) and (2) with probabilities $r$ and $s$ respectively, with $r,s\in(0,1)$. Then the lottery induced by the agent’s policy $\pi$ can be expressed as follows:
$r(a_{1}X_{1}+a_{2}X_{2}+…+a_{n}X_{n})\\ +s(b(p_{1}X_{1}+p_{2}X_{2}+…+p_{n}X_{n})\\ +(1-b)(q_{1}Y_{1}+q_{2}Y_{2}+…+q_{n}Y_{n}))+Z$
Here $a$ and $b$ denote probabilities that arise from the agent’s own stochastic choosing. Thus, $a$ and $b$ are under the agent’s control. By contrast, $p$, $q$, $r$, and $s$ are probabilities given by the environment and hence out of the agent’s control. $Z$ is a catch-all lottery that covers what happens in all situations besides those described in (1) and (2).
From the lottery induced by $\pi$, we can deduce the probabilities of each $X_{i}$, $Y_{i}$, and $X_{i}\vee Y_{i}$ given $\pi$. They are as follows:
$Pr_{\pi}\{X_{i}\}=ra_{i}+sbp_{i}$
$Pr_{\pi}\{Y_{i}\}=s(1-b)q_{i}$
$Pr_{\pi}\{X_{i}\vee Y_{i}\}=ra_{i}+sbp_{i}+s(1-b)q_{i}$
Now consider an alternative policy $\pi^{\prime}$ that makes two changes to policy $\pi$. First, the probability that the agent chooses each $X_{i}$ in (1) is modulated by a set of $\epsilon_{i}$. So in (1), the agent’s choice induces the lottery $(a_{1}+\epsilon_{1})X_{1}+(a_{2}+\epsilon_{2})X_{2}+...+(a_{n}+\epsilon_{n})X_{n}$. These $\epsilon_{i}$ are such that $\sum_{i}\epsilon_{i}=0$ and $a_{i}+\epsilon_{i}\in(0,1)$ for all $i$.
Second, the probability that the agent chooses lottery $X$ in (2) increases by $\delta$. So in (2), the agent’s choice induces the lottery $(b+\delta)(p_{1}X_{1}+p_{2}X_{2}+…+p_{n}X_{n})+(1-b-\delta)(q_{1}Y_{1}+q_{2}Y_{2}+…+q_{n}Y_{n})$.
Assume, as above, that the agent faces the situations described (1) and (2) with probabilities $r$ and $s$ respectively. Then the lottery induced by the policy $\pi^{\prime}$ can be expressed as follows:
$r((a_{1}+\epsilon_{1})X_{1}+(a_{2}+\epsilon_{2})X_{2}+...+(a_{n}+\epsilon_{n})X_{n})\\ +s((b+\delta)(p_{1}X_{1}+p_{2}X_{2}+…+p_{n}X_{n})\\ +(1-b-\delta)(q_{1}Y_{1}+q_{2}Y_{2}+…+q_{n}Y_{n}))+Z$
From the lottery induced by $\pi^{\prime}$, we can deduce the probabilities of $X_{i}$, $Y_{i}$, and $X_{i}\vee Y_{i}$ given $\pi^{\prime}$. They are as follows:
$Pr_{\pi^{\prime}}\{X_{i}\}=r(a_{i}+\epsilon_{i})+s(b+\delta)p_{i}$
$Pr_{\pi^{\prime}}\{Y_{i}\}=s(1-b-\delta)q_{i}$
$Pr_{\pi^{\prime}}\{X_{i}\vee Y_{i}\}=r(a_{i}+\epsilon_{i})+s(b+\delta)p_{i}+s(1-b-\delta)q_{i}$
We then set $Pr_{\pi}\{X_{i}\vee Y_{i}\}=Pr_{\pi^{\prime}}\{X_{i}\vee Y_{i}\}$ for each $i$ and use these equations to express each $\epsilon_{i}$ as a function of $\delta$.
$Pr_{\pi}\{X_{i}\vee Y_{i}\}=\,$ $Pr_{\pi^{\prime}}\{X_{i}\vee Y_{i}\}$
$ra_{i}+sbp_{i}+s(1-b)q_{i}=\,$ $r(a_{i}+\epsilon_{i})+s(b+\delta)p_{i}$
$+s(1-b-\delta)q_{i}$
$0=\,$ $r\epsilon_{i}+s\delta p_{i}-s\delta q_{i}$
$\epsilon_{i}=\,$ $\frac{s\delta q_{i}-s\delta p_{i}}{r}$
$\epsilon_{i}=\,$ $\frac{s\delta(q_{i}-p_{i})}{r}$
These are the values of $\epsilon_{i}$ that result in $Pr_{\pi}\{X_{i}\vee Y_{i}\}=Pr_{\pi^{\prime}}\{X_{i}\vee Y_{i}\}$.
[…]
We now prove that $\pi^{\prime}$ dominates $\pi$.
Let $c_{i}=Pr_{\pi}\{X_{i}\vee Y_{i}\}$. Let $d_{i}=Pr_{\pi}\{X_{i}|X_{i}\vee Y_{i}\}$. That lets us express the lottery induced by $\pi$ as:
$c_{1}(d_{1}X_{1}+(1-d_{1})Y_{1})+c_{2}(d_{2}X_{2}+(1-d_{2})Y_{2})\\ +…+c_{n}(d_{n}X_{n}+(1-d_{n})Y_{n})+Z$
Let $e_{i}=Pr_{\pi^{\prime}}\{X_{i}\}-Pr_{\pi}\{X_{i}\}$. That lets us express the lottery induced $\pi^{\prime}$ as:
[…]